OpenStudy (anonymous):

Please help; MEDAL WILL BE GIVEN!!! :j Find the length of an arc of 40° in a circle with an 8 inch radius.

OpenStudy (anonymous):

\[A.) 8 \pi/9 inches\]

OpenStudy (anonymous):

\[B.) 16 \pi/9 inches\]

OpenStudy (anonymous):

\[C.) 64 \pi/9 inches\]

OpenStudy (anonymous):

What is the angle of a circle? It is 360 degrees. 40/360=1/9 So the length of the arc is 1/9 of the whole perimeter of the circle.

OpenStudy (anonymous):

Now do you want to solve the question yourself or let me help you?

OpenStudy (anonymous):

I want you t help me please :j

OpenStudy (anonymous):

All right.

OpenStudy (anonymous):

First find the perimeter of the circle. \[2 \times radius=diameter\]; \[diameter \times \pi=perimeter\] So\[2 \times 8 \times \pi=16\pi\] Then \[16\pi \times \frac{ 1 }{ 9 }=\frac{ 16 }{ 9 }\pi\] B is right.

OpenStudy (anonymous):

That's what I thought but I wasn't completely sure. Thank you sooo much!!! :j

OpenStudy (anonymous):

You are welcome!